#8 AT&T 10:54 PM X Practice Chapter 4: Atte... Question 4 Find the indicated probability Refer to the table which summarizes the results of testing for a certain disease. If one of the results is randomly selected, what is the probability that it is a false positive (test indicates the person has the disease when in fact they dont)? What does this probability suggest about the accuracy of the test? 1 0.155; The probability of this error is high so the test is not very accurate. 0.414; The probability of this error ? is high so the test is not very accurate. 2 of 24 completed Save for Later Submit
media%2F778%2F77858a62-7875-4ef8-bc87-d4
Answer #1

Here we have given the probability of false positive i.e. in testing we reject H0 when it is actualy true.

Here test indicate that person has the disease but actualy that person have no disease. In testing of hypothesis such type of error is termed as type-I error.

p(Type-I error) \leqslant\alpha

We know that \alpha is the level of significance and its value depends on researcher. Generaly value of \alpha takes 0.01, 0.05 or 0.1.

Hence option (4) is correct i.e. 0.0256; the probability of this error is low so test is fairly accurate.